# Heuristic Step Planning for Learning Dynamic Bipedal Locomotion: A Comparative Study of Model-Based and Model-Free Approaches

### v1

@@ Abstract
Keywords Bipedal locomotion $\cdot$ Footstep planning $\cdot$ Reinforcement learning
@@ 1 Introduction
In many studies (Reher and Ames, 2021; Gong et al., 2019), the high-level control signal provided to the walking controller is the desired torso velocity—analogous to the control of wheeled platforms. This approach is effective on flat surfaces but lacks sufficient precision in complex terrain or constrained environments. An alternative is control through desired foot positions, which provides more reliable interaction with the environment but requires integration of a step planner.
@@ 1.1 Approaches to humanoid robot control systems
Currently, there are two main methods for controlling bipedal walking robots. These methods form the basis of most motion control approaches, with numerous hybrid strategies lying between them. The first, more traditional approach is based on classical control methods. It involves generating reference trajectories for the robot’s joints or links—either in real-time or offline. These trajectories can be obtained using full-body dynamic models, such as the Hybrid Zero Dynamics (HZD) method (Reher and Ames, 2021; Gong et al., 2019), or simplified models, such as the Linear Inverted Pendulum Model (LIPM) (Feng et al., 2016) and the Angular Momentum Linear Inverted Pendulum (ALIP) (Gong and Grizzle, 2020). Once the reference trajectories are generated, the next step is to ensure accurate tracking by the robot’s joints or links. Various controllers are used for this purpose, including: PD controllers (Gong et al., 2019), inverse dynamics controllers such as Task-Space Inverse Dynamics (TSID) (Del Prete et al., 2016), and ID-QP controllers (Shen, 2022). Trajectory generation can be considered high-level control, while tracking represents low-level control.
@@
The second approach, closely linked to the growth of modern computational power, is model-free learning, in particular Reinforcement Learning (RL). RL has shown high potential in tasks requiring adaptability and robustness, as it allows robots to learn effective control strategies through trial-and-error without precise system modeling. This approach simplifies the process of control algorithm synthesis, but it may lead to unintuitive or physically unrealistic behavior due to the agent’s freedom in exploring the entire action space. Despite being relatively new compared to classical methods, RL-based controllers have recently made significant progress, with several researchers (Siekmann et al., 2021; Liao et al., 2024) successfully implementing them on real bipedal robots.
Hybrid methods are also emerging, aiming to combine the strengths of both strategies. These approaches constrain the flexibility of RL by embedding knowledge from classical control. For instance, the HZD framework can be employed to construct a gait library of reference trajectories, while reinforcement learning is used to design a low-level controller that tracks these trajectories (Chaikovskaya et al., 2023; Li et al., 2024).
@@ 1.2 Step Planning in Locomotion Tasks
Reinforcement learning (RL) methods have shown significant potential for achieving dynamic locomotion in humanoid robots. However, many existing RL-based controllers do not explicitly consider foot placement constraints (Siekmann et al., 2021; Margolis and Agrawal, 2023; Liao et al., 2024), allowing the control system to freely adjust foot positions to maintain balance. While this approach simplifies the learning problem, it limits applicability in real-world scenarios that require precise and careful foot placement, such as walking over uneven or discontinuous surfaces. Therefore, explicitly incorporating foot placement constraints into the control system is a key requirement for ensuring stable and reliable locomotion in complex and structured environments.
@@
Recent years have seen several studies focused on integrating step planning into RL-based architectures for bipedal locomotion control. For instance, in (Wang et al., 2023), the RL policy is closely coupled with a foot placement planning module, enabling determination of feasible foot positions under complex environmental conditions. Although such architectures can enhance adaptability, they also substantially increase the complexity of the control system and computational cost. Moreover, they typically require accurate estimation of global foot positions to reliably track step placement commands at each control cycle.
@@
A similar strategy was presented in (Lee et al., 2024), where target foot positions are determined using a linear inverted pendulum model (LIPM). However, we argue that the use of classical model-based planners such as LIPM is not strictly necessary and, in fact, may introduce unnecessary complexity without significant performance gain. To test this hypothesis, we conduct a comparative analysis of two step-planning-based locomotion strategies: (i) the heuristic method proposed in this work (hereafter referred to as the Linear Step Planning, or LS, approach); and (ii) the LIPM-based approach from (Lee et al., 2024). Through this study, we aim to address a key question in RL-based bipedal locomotion: To what extent is knowledge derived from mathematical models necessary to guide the learning process? Moreover, how does the use of such models affect the behavior, efficiency, and generalization capability of the resulting control policies?
This work contributes in two main directions: (1) Integration of heuristic step planning into the learning architecture. We propose a simple step planning structure that does not require complex dynamic models (Fig. 1). Unlike previous approaches where target foot positions were considered only in the RL reward function or not defined at all (Siekmann et al., 2021; Margolis and Agrawal, 2023), our approach employs a feedback mechanism that compares actual and target foot placements during policy execution (see Fig. 1). (2) Comparative analysis of model-based methods. The proposed approach is systematically compared with a simplified dynamics model (LIPM) to assess the effect of model-derived information on learning performance, locomotion stability, motion quality, and the generalization capability of the learned policy.
The proposed approach was validated through simulations and sim-to-sim testing on the BRUCE (Bipedal Robot Unit with Compliance Enhanced) (Liu et al., 2022), demonstrating its promise for improving robotic locomotion in challenging environments. BRUCE is a kid-size humanoid robot (Fig. 2), originally developed at RoMeLa in joint effort with Westwood Robotics. It stands about 70 cm tall and weighs about 4.8 kg.
@@ 2 Preliminaries and problem formulation
A Raibert-like regulator is used to encourage the policy to accurately track a target walking velocity by adding an offset to the foot strike location (Raibert et al., 1984):
@@ 3 RL problem formulation
In an attempt to address the research question of this paper, we formulate two RL training frameworks. The first structure is shown in Fig. 1 and uses the equations presented in Section 2. In addition to this structure, we also implemented the method presented in (Lee et al., 2024).
@@
We train our policy in the Isaac Gym simulation engine on flat terrain using the PPO algorithm (Schulman et al., 2017). The state space, action space, and reward formulation for the RL problems are prepared as in (Lee et al., 2024), with the weights adjusted to fit our model as presented in Table 1. The Raibert-like regulator values used for the LS-based method are $K_{p}^{rb}=[0.3,0.3]$ and $K_{d}^{rb}=[0.1,0.1]$.
@@ 4.1.4 Energy Efficiency
The energy efficiency of the control systems was evaluated based on the mechanical and electrical costs of transportation (Torricelli et al., 2015). We simulated the robot walking on flat ground at an average speed of $0.6\,\text{m/s}$ over 100 steps. The results are presented in Table 4. The LS-based approach outperformed the LIPM-based method, even though the latter relies on robot dynamics to determine optimal positions.
@@ 4.1.5 Robustness
We evaluated the robustness of the controllers by applying external linear pushes to the robot’s base and recording its response (fall or recover). Each test consisted of 500 samples, with maximum impulse values of $7\,\text{N·s}$ and $5\,\text{N·s}$. The robot was instructed to maintain zero velocity throughout the tests.
The recovery success rates are shown in Table 5. Fig. 5 shows the distribution of coordinates where the robot either falls or successfully recovers for $I_{\text{max}}=7\,\text{N·s}$. The results show that the LS-based method outperforms the other approaches.
@@ 4.1.6 Walking on rough terrain
To evaluate the adaptability of our policy to previously unseen and uneven terrain, we conducted experiments on rough terrain (Fig. 6). The evaluation metric was the robot’s ability to maintain a predefined forward velocity command for five seconds without falling. As shown in Fig. 7, our approach achieved a higher success rate than the other approach. Although we applied the modifications proposed in (Lee et al., 2024) to improve the LIPM-based method, the simple heuristic approach exhibited approximately 60% greater stability in trials compared to the LIPM-based method. In particular, our method worked in these experiments without incorporating any information about the height of the feet into the policy.
@@ 4.1.7 Walking on a Terrain with Gaps
To evaluate the effectiveness of the proposed method, we assessed the robot’s performance on a surface containing gaps, using the same policy trained on a flat surface in the IsaacGym simulation environment. Terrain perception was based on a height map, consistent with the conditions used during training for stair traversal. When a gap was detected in the region of the next planned step, the robot shifted the foot placement to the nearest available flat surface. A similar strategy was employed in Lee et al. (2024) to enable gap crossing in legged robots.
@@ 4.2 Discussion
The LS method demonstrated an approximately 20% higher disturbance rejection capability compared to the other method under an impulse force of $5\,\text{N·s}$. Moreover, it achieved excellent results in "blind" walking over uneven terrain, providing approximately 50% higher stability in tests compared to the other method.

### v2

@@ 1 Introduction
In many studies (reher2021inverse; gong2019feedback), the high-level control signal provided to the walking controller is the desired torso velocity—analogous to the control of wheeled platforms. This approach is effective on flat surfaces but lacks sufficient precision in complex terrain or constrained environments. An alternative is control through desired foot positions, which provides more reliable interaction with the environment but requires integration of a step planner.
@@ 1.1 Approaches to humanoid robot control systems
Currently, there are two main methods for controlling bipedal walking robots. These methods form the basis of most motion control approaches, with numerous hybrid strategies lying between them. The first, more traditional approach is based on classical control methods. It involves generating reference trajectories for the robot’s joints or links—either in real-time or offline. These trajectories can be obtained using full-body dynamic models, such as the Hybrid Zero Dynamics (HZD) method (reher2021inverse; gong2019feedback), or simplified models, such as the Linear Inverted Pendulum Model (LIPM) (feng2016robust) and the Angular Momentum Linear Inverted Pendulum (ALIP) (gong2020angular). Once the reference trajectories are generated, the next step is to ensure accurate tracking by the robot’s joints or links. Various controllers are used for this purpose, including: PD controllers (gong2019feedback), inverse dynamics controllers such as Task-Space Inverse Dynamics (TSID) (delprete2016torque), and ID-QP controllers (shen2022locomotion). Trajectory generation can be considered high-level control, while tracking represents low-level control.
@@
The second approach, closely linked to the growth of modern computational power, is model-free learning, in particular Reinforcement Learning (RL). RL has shown high potential in tasks requiring adaptability and robustness, as it allows robots to learn effective control strategies through trial-and-error without precise system modeling. This approach simplifies the process of control algorithm synthesis, but it may lead to unintuitive or physically unrealistic behavior due to the agent’s freedom in exploring the entire action space. Despite being relatively new compared to classical methods, RL-based controllers have recently made significant progress, with several researchers (siekmann2021sim; liao2024berkeley) successfully implementing them on real bipedal robots.
Hybrid methods are also emerging, aiming to combine the strengths of both strategies. These approaches constrain the flexibility of RL by embedding knowledge from classical control. For instance, the HZD framework can be employed to construct a gait library of reference trajectories, while reinforcement learning is used to design a low-level controller that tracks these trajectories (chaikovskaya2023benchmarking; li2024reinforcement).
@@ 1.2 Step Planning in Locomotion Tasks
Reinforcement learning (RL) methods have shown significant potential for achieving dynamic locomotion in humanoid robots. However, many existing RL-based controllers do not explicitly consider foot placement constraints (siekmann2021sim; margolis2023walk; liao2024berkeley), allowing the control system to freely adjust foot positions to maintain balance. While this approach simplifies the learning problem, it limits applicability in real-world scenarios that require precise and careful foot placement, such as walking over uneven or discontinuous surfaces. Therefore, explicitly incorporating foot placement constraints into the control system is a key requirement for ensuring stable and reliable locomotion in complex and structured environments.
@@
Recent years have seen several studies focused on integrating step planning into RL-based architectures for bipedal locomotion control. For instance, in (wang2023learning), the RL policy is closely coupled with a foot placement planning module, enabling determination of feasible foot positions under complex environmental conditions. Although such architectures can enhance adaptability, they also substantially increase the complexity of the control system and computational cost. Moreover, they typically require accurate estimation of global foot positions to reliably track step placement commands at each control cycle.
@@
A similar strategy was presented in (lee2024integrating), where target foot positions are determined using a linear inverted pendulum model (LIPM). However, we argue that the use of classical model-based planners such as LIPM is not strictly necessary and, in fact, may introduce unnecessary complexity without significant performance gain. To test this hypothesis, we conduct a comparative analysis of two step-planning-based locomotion strategies: (i) the heuristic method proposed in this work (hereafter referred to as the Linear Step Planning, or LS, approach); and (ii) the LIPM-based approach from (lee2024integrating). Through this study, we aim to address a key question in RL-based bipedal locomotion: To what extent is knowledge derived from mathematical models necessary to guide the learning process? Moreover, how does the use of such models affect the behavior, efficiency, and generalization capability of the resulting control policies?
This work contributes in two main directions: (1) Integration of heuristic step planning into the learning architecture. We propose a simple step planning structure that does not require complex dynamic models (Fig. 1). Unlike previous approaches where target foot positions were considered only in the RL reward function or not defined at all (siekmann2021sim; margolis2023walk), our approach employs a feedback mechanism that compares actual and target foot placements during policy execution (see Fig. 1). (2) Comparative analysis of model-based methods. The proposed approach is systematically compared with a simplified dynamics model (LIPM) to assess the effect of model-derived information on learning performance, locomotion stability, motion quality, and the generalization capability of the learned policy.
The proposed approach was validated through simulations and sim-to-sim testing on the BRUCE (Bipedal Robot Unit with Compliance Enhanced) (liu2022design), demonstrating its promise for improving robotic locomotion in challenging environments. BRUCE is a kid-size humanoid robot (Fig. 2), originally developed at RoMeLa in joint effort with Westwood Robotics. It stands about 70 cm tall and weighs about 4.8 kg.
@@ 2 Preliminaries and problem formulation
A Raibert-like regulator is used to encourage the policy to accurately track a target walking velocity by adding an offset to the foot strike location (raibert1984balance):
@@ 3 RL problem formulation
In an attempt to address the research question of this paper, we formulate two RL training frameworks. The first structure is shown in Fig. 1 and uses the equations presented in Section 2. In addition to this structure, we also implemented the method presented in (lee2024integrating).
@@
We train our policy in the Isaac Gym simulation engine on flat terrain using the PPO algorithm (schulman2017ppo). The state space, action space, and reward formulation for the RL problems are prepared as in (lee2024integrating), with the weights adjusted to fit our model as presented in Table 1. The Raibert-like regulator values used for the LS-based method are $K_{p}^{rb}=[0.3,0.3]$ and $K_{d}^{rb}=[0.1,0.1]$.
@@ 4.1.4 Energy Efficiency
The energy efficiency of the control systems was evaluated based on the mechanical and electrical costs of transportation (torricelli2015benchmarking). We simulated the robot walking on flat ground at an average speed of $0.6\,\text{m/s}$ over 100 steps. The results are presented in Table 4. The LS-based approach outperformed the LIPM-based method, even though the latter relies on robot dynamics to determine optimal positions.
@@ 4.1.5 Robustness
We evaluated the robustness of the controllers by applying external linear pushes to the robot’s base and recording its response (fall or recover). Each test consisted of 500 samples, with maximum impulse values of $7\,\text{N\textperiodcentered s}$ and $5\,\text{N\textperiodcentered s}$. The robot was instructed to maintain zero velocity throughout the tests.
The recovery success rates are shown in Table 5. Fig. 5 shows the distribution of coordinates where the robot either falls or successfully recovers for $I_{\text{max}}=7\,\text{N\textperiodcentered s}$. The results show that the LS-based method outperforms the other approaches.
@@ 4.1.6 Walking on rough terrain
To evaluate the adaptability of our policy to previously unseen and uneven terrain, we conducted experiments on rough terrain (Fig. 6). The evaluation metric was the robot’s ability to maintain a predefined forward velocity command for five seconds without falling. As shown in Fig. 7, our approach achieved a higher success rate than the other approach. Although we applied the modifications proposed in (lee2024integrating) to improve the LIPM-based method, the simple heuristic approach exhibited approximately 60% greater stability in trials compared to the LIPM-based method. In particular, our method worked in these experiments without incorporating any information about the height of the feet into the policy.
@@ 4.1.7 Walking on a Terrain with Gaps
To evaluate the effectiveness of the proposed method, we assessed the robot’s performance on a surface containing gaps, using the same policy trained on a flat surface in the IsaacGym simulation environment. Terrain perception was based on a height map, consistent with the conditions used during training for stair traversal. When a gap was detected in the region of the next planned step, the robot shifted the foot placement to the nearest available flat surface. A similar strategy was employed in lee2024integrating to enable gap crossing in legged robots.
@@ 4.2 Discussion
The LS method demonstrated an approximately 20% higher disturbance rejection capability compared to the other method under an impulse force of $5\,\text{N\textperiodcentered s}$. Moreover, it achieved excellent results in ”blind” walking over uneven terrain, providing approximately 50% higher stability in tests compared to the other method.
@@ 5 CONCLUSIONS
[…]
The authors would like to express their gratitude to all colleagues and reviewers whose comments and suggestions greatly improved this work.